# Effect of Synthetic Jets Actuator Parameters on Deep Reinforcement Learning-Based Flow Control Performance in a Square Cylinder

### v1

@@ Abstract
We utilize deep reinforcement learning (DRL) algorithms to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of a square cylinder, achieving effective active flow control. Through DRL-based active flow control (AFC) technology, we significantly reduce the lift and drag coefficients of the square cylinder at $Re=100$ and $Re=500$, while completely suppressing vortex shedding in the wake flow field. Additionally, we conduct a sensitivity analysis of the position and width parameters of the synthetic jets regarding flow control performance. Our observations indicate that positioning the synthetic jets near the trailing edge corners of the square cylinder, as opposed to the leading edge corners, completely suppresses vortex shedding, resulting in more stable controlled flow coefficients, $C_{D}$ and $C_{L}$. When the synthetic jets are positioned at the trailing edge corners, flow control reduces the mean drag coefficient by 14.4% and the standard deviation of the lift coefficient by 86.1% for the baseline flow at $Re=100$. For the baseline flow at $Re=500$, flow control reduces the mean drag coefficient by 51.4% and the standard deviation of the lift coefficient by 90.5%. At both Reynolds numbers, vortex shedding in the wake flow field is completely suppressed. Furthermore, using narrower synthetic jets results in a lower reduction rate of the standard deviations of $C_{D}$ and $C_{L}$, while increasing the mean and standard deviation of the mass flow rate of the jets used for flow control. This study provides guidance on optimizing the width and position of synthetic jets for DRL-based active flow control.
@@ I INTRODUCTION
Given the significant potential of DRL in control and optimization, a large number of scholars have been applying it in the field of flow control. Rabault et al.Rabault et al. (2019) pioneered the application of DRL to AFC technologies, achieving an 8% reduction in drag around a cylinder at $Re=100$. This seminal work sparked interest in applying DRL techniques within the field of fluid dynamics. Tang et al.Tang et al. (2020a) focused on the robustness across different Reynolds numbers, while Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021) explored the impact of probe distribution on control performance. Li and ZhangLi and Zhang (2022) applied DRL-based AFC to cylinders under various blockage ratios, integrating flow physics mechanisms during the control process to achieve significant outcomes. Fan et al.Fan et al. (2020) developed an active flow control strategy based on DRL through experimental fluid dynamics, demonstrating the effectiveness of applying reinforcement learning in experimental fluid mechanics. Rabault and KuhnleRabault and Kuhnle (2019a) and Jia and XuJia and Xu (2024a) addressed issues related to parallel strategies and training when coupling Computational Fluid Dynamics (CFD) codes with DRL framework in parallel computing environments. Wang et al.Wang et al. (2022) developed an open source library that integrated CFD code OpenFoam with DRL framework TensorForce. DRL has demonstrated excellent control performance in the field of AFC, sparking increasing exploration. While there is a plethora of related work, we will not delve into each individually. Instead, we summarize in Table. 1 the research on the application of DRL-based AFC methods in the context of flow around bluff bodies.
@@
Re Bluff Body Reference Strategy Control Algorithm Solver Drag Reduction Vortex Suppressed Contribution
100 Cylinder Rabault et al.Rabault et al. (2019) Synthetic Jets PPO FeniCS 8% - Pioneer
100 Cylinder Wang et al.Wang et al. (2022) Synthetic Jets PPO OpenFOAM 8% - Python library
100 Cylinder Castellanos et al.Castellanos et al. (2022) Synthetic Jets LGPC/PPO FeniCS 8% - DRL and LGPC
100 Cylinder Jia and XuJia and Xu (2024b) Synthetic Jets PPO OpenFOAM 8% YES Parallel strategy
100 Cylinder Ren, Wang, and TangRen, Wang, and Tang (2021) WSLB PPO LBM - - Fast training speed
@@
120 Cylinder Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021) Synthetic Jets S-PPO-CMA FastS 18.4% - Sensor placement
$\leq$200 Cylinder Li and ZhangLi and Zhang (2022) Synthetic Jets PPO Nek5000 - YES Embedding physical
100-300 Cylinder He et al.He et al. (2023) Synthetic Jets PPO OpenFOAM 6%-24% - Transfer learning
100-400 Cylinder Tang et al.Tang et al. (2020b) Synthetic Jets PPO FeniCS 6%-39% - Robustness for Re
1,000 Cylinder Ren, Rabault, and TangRen, Rabault, and Tang (2021) Synthetic Jets PPO LBM 30% - Weakly turbulent
10,160 Cylinder Fan et al.Fan et al. (2020) Cylinders TD3 Lilypad 30% - Combine experiment
100 Ellipse Jia and XuJia and Xu (2024b) Synthetic Jets PPO OpenFOAM 16% YES Sensitivity of $Ar$
100 Square Wang et al.Wang et al. (2022) Synthetic Jets SAC OpenFOAM 14% - Python library
100 Square Chen et al.Chen et al. (2023) Synthetic Jets SAC OpenFOAM 14% YES Jet position
100 Square Xia et al.Xia et al. (2024) Synthetic Jets SAC and TQC FEniCS 17.3% - Partial measurements
100-400 Square Jia and XuJia and Xu (2024c) Synthetic Jets SAC OpenFOAM 14%-47% YES Multi-objective
500-2,000 Square Yan et al.Yan et al. (2023) Synthetic Jets SAC OpenFOAM 44%-61% - Higher Re
The research in the above table jointly promotes the application of DRL in the field of active flow control. It encompasses developments from low to relatively high Reynolds numbers, transitions from 2D to 3D configurations, shifts from numerical simulations to fluid experiments, and geometric changes from circular to elliptical bodies. Additionally, it includes advancements in DRL algorithms, software development, and hardware progress. Given the success of DRL-based AFC techniques in effectively managing flow around cylinders, an increasing number of researchers have contributed to this field. Their research significantly extend beyond the initial study of flows past a cylinder at $Re=100$ by Rabault et al.Rabault et al. (2019). This results in significant pressure fluctuations and a higher drag coefficient on the back side of the square cylinder compared to cylinder shapes, due to more pronounced fluid separation and broader wake regions.Yoon, Yang, and Choi (2010); Sen, Mittal, and Biswas (2011) The unique flow characteristics of flow around square cylinders have garnered significant interest among researchers.
@@
Wang et al.Wang et al. (2022) initially applied the Soft Actor-Critic (SAC) algorithm to active flow control of a square cylinder, achieving a 13.7% reduction in drag at $Re=100$. Despite this significant reduction, vortex shedding behind the cylinder is still evident. Xia et al.Xia et al. (2024) conduct active flow control around a square cylinder at $Re=100$ using RL, innovatively training with an augmented state that includes both current and past measurements and actions, which is represented as a nonlinear autoregressive exogenous model to address issues of partial observability. Jia and XuJia and Xu (2024c) employed the SAC algorithm to regulate the flow from synthetic jets positioned on the upper and lower surfaces of a constrained square cylinder at $Re=100$, 200, 300, and 400. The research results demonstrated that the active control strategy based on the SAC algorithm effectively reduces drag by approximately 14.4%, 26.4%, 38.9%, and 47.0% respectively. A distinctive feature of this study compared to others is its success in suppressing vortex shedding behind the square cylinder across the range of Reynolds numbers from 100 to 400. This DRL-based AFC delivers an energy-efficient control strategy that not only reduces drag and suppresses lift but also effectively suppresses vortex shedding. Yan et al.Yan et al. (2024) investigated the impact of multiple jet actuators on drag and lift reduction for rectangular cylinders with four different aspect ratios at $Re=1,000$, employing DRL for control. Specifically, the deployment of four synthetic jets resulted in an average drag reduction of 63.2%, while eight jets achieved a 77.1% reduction. This study investigated various jet distribution strategies, emphasizing the importance of multi-jet actuation in enhancing aerodynamic performance.
@@
Our research is motivated by two main reasons. Firstly, there is currently no study that has successfully used AFC to achieve both efficient drag reduction and complete suppression of vortex shedding at Reynolds numbers greater than 500. Secondly, previous studies have shown that the effects of synthetic jets placement on control performance exhibit different outcomes at $Re=100$ and $Re=500$. We are interested in exploring whether it is possible, at $Re=100$ and $Re=500$, to use AFC based on DRL to achieve both efficient drag reduction and complete suppression of vortex shedding. Additionally, we design two synthetic jet placements and four synthetic jet widths to conduct a thorough analysis of their impacts on flow control performance. This analysis will provide valuable references for the design of synthetic jets. The structure of this paper is as follows. In Sec. II, we introduce the DRL framework, which consists of a fluid simulation environment and the SAC agent. In Sec. III, we present and discuss the results. Initially, we describe the Baseline and the Base flow. Then, we discuss the convergence performance of DRL at $Re=100$ and $Re=500$. Finally, we analyze the effects of the placement and width of synthetic jets on control performance. In Sec. IV, we summarize the main conclusions of this study and discuss the broader implications of these findings.
@@ II.1 Flow environment
$Re=\frac{\overline{U}D}{\nu}$ is the Reynolds number, where $\overline{U}$ is the mean velocity at the inlet, $D$ is the diameter of the cylinder and $\nu$ is the kinematic viscosity. Together, these equations articulate the fundamental behaviors of fluid flow, linking the dynamics of velocity changes to the principle of mass conservation in fluid mechanics.
@@
To verify the accuracy of the grid, this study conducted a grid convergence investigation. We compared results from three different grid scales and contrasted them with computational outcomes from other studies, summarizing the findings in Table. 2. The table lists the values for the mean drag coefficient ($C_{D,Mean}$), the standard deviation of the drag coefficient ($C_{D,Std}$), the standard deviation of the lift coefficient ($C_{L,Std}$), and the Strouhal number ($S_{t}$) at $Re=100$ and $Re=500$. All three grids utilized structured meshes, differing only in the level of refinement, with the number of grid cells detailed in the table. The computational results in Table. 2 demonstrate that the differences among the three grid resolutions are insignificant and consistent with values obtained in existing literature. In the subsequent advanced DRL training phase, the main grid scheme is employed to balance computational accuracy and cost-effectiveness, ensuring the required precision is achieved while minimizing computational overhead.
@@ III.1 Baseline flow simulation
At $Re=100$, the wake region of the square cylinder exhibits regular vortex shedding, as shown in Fig. 8. Vortices shed from the trailing edge of the cylinder form a regular pattern, characterized by alternating vortices along the centerline, creating a street-like structure. At $Re=500$, the shedding pattern in the wake changes, as depicted in Fig. 9. After flow separation occurs at the leading edge of the square cylinder, some separated fluid reattaches to the cylinder surface. The vortex shedding in the wake region becomes increasingly unstable, with paired and single vortices shedding together, exhibiting a $P+S$ wake mode. Compared to $Re=100$, the flow instability at $Re=500$ intensifies, leading to a more complex shedding pattern in the wake. We have described the flow structures and vortex shedding patterns of the baseline flow at $Re=100$ and $Re=500$. At Reynolds number 500, the flow exhibits greater instability and complexity, indicating that the difficulty of flow control is increased. Implementing flow control techniques in such conditions requires advanced strategies that can effectively manipulate the flow dynamics and mitigate flow separation.
@@ III.3 Sensitivity analyses of synthetic jets
  • a. $C_{D,Baseline}$ represents the time-averaged drag coefficient of the square cylinder once the baseline flow has fully developed. $C_{D,DRL}$ represents the time-averaged drag coefficient of the square cylinder after AFC based on DRL.
  • b. The Reduction is quantified and presented in percentage (%).
  • c. Action represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the inlet.
@@ IV CONCLUSIONS
In this study, we investigate the application of DRL-based AFC in mitigating the lift and drag coefficients of a square cylinder while suppressing vortex shedding in the wake field. We design an interaction framework between the agent and a numerical simulation environment, optimizing its control strategy through learning and trial-and-error. The agent observes the state information of the numerical simulation environment and adjusts the mass flow rates of two synthetic jets on the square cylinder based on feedback from reward signals. Through an iterative optimization process, the SAC agent gradually converges towards the optimal flow control strategy, achieving multiple control objectives such as reducing lift and drag coefficients and suppressing vortex shedding. The primary objective of this work is to evaluate the performance of the SAC agent in controlling the synthetic jets on the square cylinder for active flow control, and to further analyze the influence of the position and width of the synthetic jets actuators on flow control effectiveness. The main findings of this study can be summarized as follows:

### v2

@@ Abstract
We utilize deep reinforcement learning (DRL) algorithms to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of a square cylinder for active flow control. Through DRL-based active flow control (AFC) technology, we significantly reduce the lift and drag coefficients of the square cylinder at Reynolds number ($Re$) $=100$ and $Re=500$, while completely suppressing vortex shedding in the wake flow field. Additionally, we conduct a sensitivity analysis of the position and width parameters of the synthetic jets regarding flow control performance. Our observations indicate that positioning the synthetic jets near the trailing edge corners of the square cylinder, rather than the leading edge corners, can completely suppress vortex shedding, resulting in more stable lift and drag coefficients in the controlled flow. When the synthetic jets are positioned at the trailing edge corners, flow control reduces the mean drag coefficient by 14.4% and the standard deviation of the lift coefficient by 86.1% for the baseline flow at $Re=100$. For the baseline flow at $Re=500$, flow control reduces the mean drag coefficient by 51.4% and the standard deviation of the lift coefficient by 90.5%. At both Reynolds numbers, vortex shedding in the wake flow field is completely suppressed. Furthermore, using narrower synthetic jets results in a lower reduction rate of the standard deviations of the lift and drag coefficients, while increasing the mean and standard deviation of the mass flow rate of the jets used for flow control. This study provides guidance on optimizing the width and position of synthetic jets for DRL-based active flow control.
@@ I INTRODUCTION
Given the significant potential of DRL in control and optimization, a large number of scholars have been applying it in the field of flow control. Rabault et al.Rabault et al. (2019) pioneered the application of DRL to AFC technologies, achieving an 8% reduction in drag around a cylinder at $Re=100$. This seminal work sparked interest in applying DRL techniques within the field of fluid dynamics. Tang et al.Tang et al. (2020a) focused on the robustness across different Reynolds numbers, while Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021) explored the impact of probe distribution on control performance. Li and ZhangLi and Zhang (2022) applied DRL-based AFC to cylinders under various blockage ratios, integrating flow physics mechanisms during the control process to achieve significant outcomes. Fan et al.Fan et al. (2020) developed an active flow control strategy based on DRL through experimental fluid dynamics, demonstrating the effectiveness of applying reinforcement learning in experimental fluid mechanics. Rabault and KuhnleRabault and Kuhnle (2019a) and Jia and XuJia and Xu (2024a) addressed issues related to parallel strategies and training when coupling Computational Fluid Dynamics (CFD) codes with DRL framework in parallel computing environments. Wang et al.Wang et al. (2022) developed an open source library that integrated CFD code OpenFoam with DRL framework TensorForce. DRL has demonstrated excellent control performance in the field of AFC, sparking increasing exploration. There has been applications from low to relatively high Reynolds numbers, transitions from 2D to 3D configurations, shifts from numerical simulations to fluid experiments, and geometric changes in the body of interest from circular to square bodies. While there is a plethora of related work, we will not delve into each individually. Instead, we summarize in Table. 1 some relevant research on the application of DRL-based AFC methods in the context of flow around bluff bodies.
@@
Re Bluff Body Reference Strategy Control Algorithm Solver Drag Reduction Vortex Suppressed
100 Cylinder Rabault et al.Rabault et al. (2019) Synthetic Jets PPO FeniCS 8% -
100 Cylinder Wang et al.Wang et al. (2022) Synthetic Jets PPO OpenFOAM 8% -
100 Cylinder Castellanos et al.Castellanos et al. (2022) Synthetic Jets LGPC/PPO FeniCS 8% -
100 Cylinder Jia and XuJia and Xu (2024b) Synthetic Jets PPO OpenFOAM 8% YES
100 Cylinder Ren, Wang, and TangRen, Wang, and Tang (2021) WSLB PPO LBM - -
@@
120 Cylinder Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021) Synthetic Jets S-PPO-CMA FastS 18.4% -
$\leq$200 Cylinder Li and ZhangLi and Zhang (2022) Synthetic Jets PPO Nek5000 - YES
100-300 Cylinder He et al.He et al. (2023) Synthetic Jets PPO OpenFOAM 6%-24% -
100-400 Cylinder Tang et al.Tang et al. (2020b) Synthetic Jets PPO FeniCS 6%-39% -
1,000 Cylinder Ren, Rabault, and TangRen, Rabault, and Tang (2021) Synthetic Jets PPO LBM 30% -
10,160 Cylinder Fan et al.Fan et al. (2020) Cylinders TD3 Lilypad 30% -
100 Ellipse Jia and XuJia and Xu (2024b) Synthetic Jets PPO OpenFOAM 16% YES
100 Square Wang et al.Wang et al. (2022) Synthetic Jets SAC OpenFOAM 14% -
100 Square Chen et al.Chen et al. (2023) Synthetic Jets SAC OpenFOAM 14% YES
100 Square Xia et al.Xia et al. (2024) Synthetic Jets SAC and TQC FEniCS 17.3% -
100-400 Square Jia and XuJia and Xu (2024c) Synthetic Jets SAC OpenFOAM 14%-47% YES
500-2,000 Square Yan et al.Yan et al. (2023) Synthetic Jets SAC OpenFOAM 44%-61% -
The research in the above table demonstrate the wide range of applications of DRL in the field of active flow control. These research significantly extend beyond the initial study of flow past a cylinder at $Re=100$ by Rabault et al.Rabault et al. (2019). However, the effectiveness of DRL control algorithm varies across different work. This is especially evident among the body of research on square cylinders. As the sharp corners and straight edges force the fluid to separate at the leading edge, flows around square cylinders are characterized by complex wake structures and shedding vortices.Bai and Alam (2018a); Yoon, Yang, and Choi (2010); Sen, Mittal, and Biswas (2011) The highly nonlinear and unique flow features makes it extremely challenging to determine the optimal control strategy. Hence, many different control experiments have been devised heuristically, leading to varying degrees of control effectiveness.
@@
Wang et al.Wang et al. (2022) initially applied the Soft Actor-Critic (SAC) algorithm to active flow control of a square cylinder, achieving a 13.7% reduction in drag at $Re=100$. Despite this significant reduction, vortex shedding behind the cylinder is still evident. Xia et al.Xia et al. (2024) conduct active flow control around a square cylinder at $Re=100$ using RL, training with an augmented state that includes both current and past measurements and actions, which are represented by a nonlinear auto-regressive exogenous model to address issues of partial observability. Jia and XuJia and Xu (2024c) employed the SAC algorithm to regulate the flow from synthetic jets positioned on the upper and lower surfaces of a constrained square cylinder at $Re=100$, 200, 300, and 400. The research results demonstrated that the active control strategy based on the SAC algorithm effectively reduces drag by approximately 14.4%, 26.4%, 38.9%, and 47.0% respectively. A distinctive feature of this study compared to others is its success in suppressing vortex shedding behind the square cylinder across the range of Reynolds numbers from 100 to 400. This DRL-based AFC delivers an energy-efficient control strategy that not only reduces drag and suppresses lift but also effectively suppresses vortex shedding. Yan et al.Yan et al. (2024) investigated the impact of multiple jet actuators on drag and lift reduction for rectangular cylinders with four different aspect ratios at $Re=1,000$, employing DRL for control. Specifically, the deployment of four synthetic jets resulted in an average drag reduction of 63.2%, while eight jets achieved a 77.1% reduction. This study investigated various jet distribution strategies, emphasizing the importance of multi-jet actuation in enhancing aerodynamic performance.
@@
Given the current state of the art in DRL-based AFC of flow around square cylinders, it is evident that the there is a lack of parameterization for synthetic jets to reconcile the varying degrees of control effectiveness in literature. Consequently, our research is motivated by two main reasons. Firstly, there is currently no study that has successfully used AFC to achieve both efficient drag reduction and complete suppression of vortex shedding at Reynolds numbers greater than 500. Secondly, previous studies have shown that the effects of synthetic jets placement on control performance exhibit different outcomes at $Re=100$ and $Re=500$. We are interested in exploring whether it is possible, at $Re=100$ and $Re=500$, to use AFC based on DRL to achieve both efficient drag reduction and complete suppression of vortex shedding. Additionally, we design two synthetic jet placements and four synthetic jet widths to conduct a thorough analysis of their impacts on flow control performance. The parameterized design of control experiments and corresponding analysis aim to provide valuable references for the control of flow around square cylinders. The structure of this paper is as follows. In Sec. II, we introduce the DRL framework, which consists of a fluid simulation environment and the SAC agent. In Sec. III, we present and discuss the results. Initially, we describe the Baseline and the Base flow. Then, we discuss the convergence performance of DRL at $Re=100$ and $Re=500$. Finally, we analyze the effects of the placement and width of synthetic jets on control performance. In Sec. IV, we summarize the main conclusions of this study and discuss the broader implications of these findings.
@@ II.1 Flow environment
Here, $Re=\frac{\overline{U}D}{\nu}$ is the Reynolds number, where $\overline{U}$ is the mean velocity at the inlet and $\nu$ is the kinematic viscosity. Together, these equations articulate the fundamental behaviors of fluid flow, linking the dynamics of velocity changes to the principle of mass conservation in fluid mechanics.
@@
To verify the accuracy of the grid, this study conducted a grid convergence investigation. We compared results from three different grid scales and contrasted them with computational outcomes from other studies, summarizing the findings in Table. 2. The table lists the values for the mean drag coefficient ($C_{D,Mean}$), the standard deviation of the drag coefficient ($C_{D,Std}$), the standard deviation of the lift coefficient ($C_{L,Std}$), and the $S_{t}$ at $Re=100$ and $Re=500$. All three grids utilized structured meshes, differing only in the level of refinement, with the number of grid cells detailed in the table. The computational results in Table. 2 demonstrate that the differences among the three grid resolutions are insignificant and consistent with values obtained in existing literature. In the subsequent advanced DRL training phase, the main grid scheme is employed to balance computational accuracy and cost-effectiveness, ensuring the required precision is achieved while minimizing computational overhead.
@@ III.1 Baseline flow simulation
At $Re=100$, the wake region of the square cylinder exhibits regular vortex shedding, as shown in Fig. 8. Vortices shed from the trailing edge of the cylinder form a regular pattern, characterized by alternating vortices along the centerline, creating a street-like structure. At $Re=500$, the shedding pattern in the wake changes, as depicted in Fig. 9. After flow separation occurs at the leading edge of the square cylinder, some separated fluid reattaches to the cylinder surface. The vortex shedding in the wake region of the square cylinder becomes increasingly unstable, with vortices squeezing, stretching, and deforming as they develop downstream in the form of vortex pairs. Compared to $Re=100$, the flow instability at $Re=500$ intensifies, leading to a more complex shedding pattern in the wake. We have described the flow structures and vortex shedding patterns of the baseline flow at $Re=100$ and $Re=500$. At Reynolds number 500, the flow exhibits greater instability and complexity, indicating that the difficulty of flow control is increased. Implementing flow control techniques in such conditions requires advanced strategies that can effectively manipulate the flow dynamics and mitigate flow separation.
@@ III.3 Sensitivity analyses of synthetic jets
  • a. $C_{D,Baseline}$ represents the time-averaged drag coefficient of the square cylinder when the baseline flow is fully developed.
  • b. $C_{D,DRL}$ represents the time-averaged drag coefficient of the square cylinder after AFC based on DRL.
  • c. The reduction is used to quantify the reduction ratio of the DRL control result relative to the baseline, expressed in percentage (%).
  • d. Action represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the inlet.
@@ IV CONCLUSIONS
In this study, we investigate the application of DRL-based AFC in mitigating the lift and drag coefficients of a square cylinder while suppressing vortex shedding in the wake field. We design an interaction framework between the agent and a numerical simulation environment, optimizing its control strategy through learning and trial-and-error. The agent observes the state information of the numerical simulation environment and adjusts the mass flow rates of two synthetic jets on the square cylinder based on feedback from reward signals. Through an iterative optimization process, the SAC agent gradually converges towards the optimal flow control strategy, achieving multiple control objectives such as reducing lift and drag coefficients and suppressing vortex shedding. We will evaluate the performance of the SAC agent in controlling synthetic jets on a square cylinder for active flow control, and further analyze the impact of the position and width of the synthetic jet actuators on flow control effectiveness. The main findings of this study can be summarized as follows: